# Evaluating Memory in LLM Agents via Incremental Multi-Turn Interactions

### v2

@@ 4.1 Experimental Setup
The datasets are split into four categories and the statistics of all datasets are also shown in Table 5. The evaluation metrics for all datasets are shown in Table 1, along with more dataset details. For the agents, as described in Section 3.2, we consider three categories of agents: Long-Context Agents, RAG agents and Agentic Memory Agents, where RAG Agents can be further split into Simple RAG Agents, Embedding-based RAG Agents and Structure-Augmented RAG Agents. We give the detailed introduction of each memory agent in Appendix B. For chunk size settings, we choose a chunk size of 512 for the SH-Doc QA, MH-Doc QA, and LME(S*) tasks in AR, as well as for all tasks in SF. This is mainly because these tasks are composed of long texts synthesized from multiple short texts. For other tasks, we use a chunk size of 4096. Considering computational overhead and API cost, we uniformly use a chunk size of 4096 for the Mem0, Cognee, Zep, and MIRIX. We report the settings of the chunk size in Table 14 in Appendix D.
@@ 4.3 Analysis and Ablation Study
In this section, we present experiments and analysis along five dimensions: input chunk size, retrieval top-, backbone model, dataset validation, and computational latency. Additional results are provided in the appendix, including context length analysis (Appendix D.4), latency and GPU memory usage comparisons (Appendix D.5, D.6), as well as further details on chunk size and top- ablations (Appendix D.2, D.3).
@@ 4.3.2 Ablation Study on Retrieval TopK
In our experiments, although we report most results with the number of retrieved chunks set to 10 in Table 4.1, we also conducted ablation studies with varying retrieval sizes. A subset of these results is visualized in Figure 3, with the full results provided in Table 9 in Appendix D. The results indicate that increasing the number of retrieved chunks generally improves performance across most tasks. It is worth noting that, with a chunk size of 4096 tokens, retrieving 10 chunks already yields an input of approximately 40k tokens. This places significant demands on model capacity. Due to this high token volume, we do not evaluate settings with 20 retrieved chunks.
@@ 5 Conclusion
Appendix A Details of Dataset
@@
A.1 Accurate Retrieval (AR)
A.1.1 Definition of AR
@@ A.1.2 Details on AR datasets
A.2 Test-time Learning (TTL)
A.2.1 Definition of TTL
@@ A.2.2 Details on TTL datasets
A.3 Long-Range Understanding (LRU)
A.3.1 Definition of LRU
@@ A.3.2 Details on LRU datasets
A.4 Selective Forgetting (SF)
A.4.1 Definition of SF
@@
A.4.2 Details on SF datasets
We use counterfactual edit pairs from the MQUAKE (Zhong et al., 2023) dataset. Each sentence containing information was assigned a number. For each edit pair, the sentence representing outdated information (the distractor) is given a smaller number, while the sentence representing more recent information (the one containing the answer) is given a larger number. We then concatenate these sentences into a long context in order according to their assigned numbers. We evaluate the SF via two datasets: Single-Hop FactConsolidation and Multi-Hop FactConsolidation. In these tasks, the agent’s responses are mostly informational entities. Therefore, we also use SubEM (Substring Exact Match) as the evaluation metric to calculate the accuracy of QA.
A.5 Justification for competencies based on cognitive science
@@
Beyond the acquisition–consolidation axis, another equally fundamental challenge is selective forgetting. Overlapping or contradictory traces can impede retrieval, and interference has long been recognized as a primary driver of forgetting in cognitive psychology (Anderson & Neely, 1996). Neurocognitive evidence further shows that the brain engages targeted control mechanisms to resolve such interference at retrieval time (Wimber et al., 2015). We therefore include selective forgetting—the ability to handle interference and contradictions—as a core dimension.
@@
Appendix B Detailed Memory Agents Description
@@
B.1 Long-Context Agents
We evaluate five modern long-context LLMs: GPT-4o (OpenAI, 2025b) serves as the high-performance, low-latency model with better cost efficiency than prior generations. While GPT-4o-mini is a lightweight, budget-friendly alternative that enables large-scale evaluations by delivering faster responses and lower per-token costs. Notably, the GPT-4.1 (OpenAI, 2025a) family strengthens instruction following and maintains strong performance at very large context windows (reported up to 1M tokens). Considering the higher token cost, we choose the GPT-4.1-mini in evaluation. Gemini-2.0-Flash (DeepMind, 2025) targets high throughput and the use of built-in tools, offering a 1M token context window for efficient long-context processing. Claude-3.7-Sonnet (Anthropic, 2025) is a hybrid-reasoning model with optional visible “extended thinking,” strong math/coding skills, and developer-controlled thinking budgets.
B.2 RAG Agents
@@
We implement a BM25 (Robertson & Walker, 1994) retriever as a strong lexical baseline: it scores documents by term frequency with saturation and inverse document frequency, with length normalization controlled by parameters and . BM25 remains competitive for exact-match queries and complements dense retrievers with robust precision on keyworded questions.
@@
B.3 Agentic Memory Agents
For Agentic Memory Agents, We evaluate the Self-RAG (Asai et al., 2023), MemGPT (Packer et al., 2023), and MIRIX (Wang & Chen, 2025) on our benchmark. Self-RAG use LLMs as the agent to decide when/what to retrieve and to critique its own outputs. MemGPT operates the hierarchical memory management, paging relevant snippets between short-term and long-term stores and using event-driven interrupts to maintain coherence and evolvability over extended interactions. MIRIX adopts a multi-agent memory architecture with six specialized memory types (Core, Episodic, Semantic, Procedural, Resource, Knowledge Vault) and a coordinator that orchestrates updates/retrieval across agents.
@@ Appendix C Prompts
C.1 Instructions for Memory Construction
@@
C.2 Instructions for Task Execution
@@
Appendix D Detailed Experimental Results
@@
D.1 Detailed Results on TTL
We give detailed results on Multi-Class Classification (MCC) task in Table D. For all three types of tasks, RAG-based agents generally underperform compared to their respective GPT-4o-mini backbones. This observation highlights certain limitations inherent to the RAG approach. For instance, in TTL tasks, RAG-based methods often struggle to more accurately retrieve context from memory that is closely associated with the input.
D.2 Results on Input Chunk Size Ablation Study
@@
D.3 Results on Retrieval TopK Ablation Study
@@
D.4 Results on Different Context Length Ablation Study
@@
D.5 Results on Computational Latency Analysis
To illustrate the latency of various memory agents in terms of (1) Memory Construction (M.C.); (2) Query Execution (Q.E.), we report the latency of various memory agents on MH-QA and LME (S*). This part of experiments is done on a server with Four NVDIA L40 GPU and AMD EPYC 7713 64-Core CPU. We use the NV-Embed-v2 (7B) as the embedding model in HippoRAG-v2. We show the results in Table 11 and 11. From the table, we find that using a smaller chunk size requires significantly more time for memory construction, especially for methods such as HippoRAG-v2, Mem0, Cognee, and MemGPT. Meanwhile, methods such as Mem0, Cognee and MIRIX need extremely high resources when constructing the memory.
@@ D.5 Results on Computational Latency Analysis
D.6 GPU Memory Usage Comparison
@@
Appendix E Experimental Settings
@@ E.2 Settings of the RAG Agents
E.3 Settings of the Chunk Size

### v3

@@ 4.1 Experimental Setup
The datasets are split into four categories and the statistics of all datasets are also shown in Table 6. The evaluation metrics for all datasets are shown in Table 2, along with more dataset details. For the agents, as described in Section 3.2, we consider three categories of agents: Long-Context Agents, RAG agents and Agentic Memory Agents, where RAG Agents can be further split into Simple RAG Agents, Embedding-based RAG Agents and Structure-Augmented RAG Agents. We give the detailed introduction of each memory agent in Appendix C. For chunk size settings, we choose a chunk size of 512 for the SH-Doc QA, MH-Doc QA, and LME(S*) tasks in AR, as well as for all tasks in SF. This is mainly because these tasks are composed of long texts synthesized from multiple short texts. For other tasks, we use a chunk size of 4096. Considering computational overhead and API cost, we uniformly use a chunk size of 4096 for Mem0, Cognee, Zep, and MIRIX. We report the detailed settings of the chunk size in Table 15 in Appendix E.
@@ 4.3 Analysis and Ablation Study
In this section, we present experiments and analysis along five dimensions: input chunk size, retrieval top-, backbone model, and dataset validation. Additional results are provided in the appendix, including compuational latency (Appendix E.5), context length analysis (Appendix E.4), latency and GPU memory usage comparisons (Appendix E.5, E.6), further details on chunk size and top- ablations (Appendix E.2, E.3), as well as the Cost-Performance estimation (Appendix I).
@@ 4.3.2 Ablation Study on Retrieval TopK
In our experiments, although we report most results with the number of retrieved chunks set to 10 in Table 4.1, we also conducted ablation studies with varying retrieval sizes. A subset of these results is visualized in Figure 3, with the full results provided in Table 10 in Appendix E. The results indicate that increasing the number of retrieved chunks generally improves performance across most tasks. It is worth noting that, with a chunk size of 4096 tokens, retrieving 10 chunks already yields an input of approximately 40k tokens. This places significant demands on model capacity. Due to this high token volume, we do not evaluate settings with 20 retrieved chunks.
@@ 5 Conclusion
Ethics Statement
This work adheres to the ICLR Code of Ethics and associated author guidance; we assess potential impacts and document mitigations accordingly. We evaluate memory in LLM agents using dialogs , license-compliant corpora; no personally identifiable information or data from minors were collected. To reduce dual-use risks, we release only safety-screened prompts and provide usage notes discouraging surveillance-oriented applications. We will release code under the MIT License and datasets/benchmark artifacts under CC BY 4.0; third-party materials retain their original licenses.
Reproducibility Statement
Upon acceptance, we will open-source all code and data used in this paper. The repository will include (i) training/evaluation scripts, configuration files, and exact prompts; (ii) dataset releases and generation scripts with seeds to fully regenerate interactions; and (iii) end-to-end run recipes. We will pin software dependencies and provide a containerized environment (Dockerfile plus conda/requirements.txt) and report hardware, CUDA/cuDNN, and OS details to support deterministic re-runs, in line with community guidance on reproducibility statements and artifact preparation.
Appendix A The Use of Large Language Models (LLMs)
In this paper writing process, we used an LLM to assist with content polishing—for example, identifying grammatical errors and suggesting revisions for sentences that were unclear or potentially ambiguous. Additionally, we used the LLM to generate character icons, which were later used in the creation of our main plot visualization.
Appendix B Details of Dataset
@@
B.1 Accurate Retrieval (AR)
B.1.1 Definition of AR
@@ B.1.2 Details on AR datasets
B.2 Test-time Learning (TTL)
B.2.1 Definition of TTL
@@ B.2.2 Details on TTL datasets
B.3 Long-Range Understanding (LRU)
B.3.1 Definition of LRU
@@ B.3.2 Details on LRU datasets
B.4 Selective Forgetting (SF)
B.4.1 Definition of SF
@@
B.4.2 Details on SF datasets
We use counterfactual edit pairs from the MQUAKE (Zhong et al., 2023b) dataset. Each sentence containing information was assigned a number. For each edit pair, the sentence representing outdated information (the distractor) is given a smaller number, while the sentence representing more recent information (the one containing the answer) is given a larger number. We then concatenate these sentences into a long context in order according to their assigned numbers. We evaluate the SF via two datasets: Single-Hop FactConsolidation and Multi-Hop FactConsolidation. In these tasks, the agent’s responses are mostly informational entities. Therefore, we also use SubEM (Substring Exact Match) as the evaluation metric to calculate the accuracy of QA.
B.5 Justification for competencies based on cognitive science
@@
Beyond the acquisition–consolidation axis, another equally fundamental challenge is selective forgetting. Overlapping or contradictory traces can impede retrieval, and interference has long been recognized as a primary driver of forgetting in cognitive psychology (Anderson and Neely, 1996). Neurocognitive evidence further shows that the brain engages targeted control mechanisms to resolve such interference at retrieval time (Wimber et al., 2015). We therefore include selective forgetting—the ability to handle interference and contradictions—as a core dimension.
@@
Appendix C Detailed Memory Agents Description
@@
C.1 Long-Context Agents
We evaluate five modern long-context LLMs: GPT-4o (OpenAI, 2025a) serves as the high-performance, low-latency model with better cost efficiency than prior generations. While GPT-4o-mini is a lightweight, budget-friendly alternative that enables large-scale evaluations by delivering faster responses and lower per-token costs. Notably, the GPT-4.1 (OpenAI, 2025b) family strengthens instruction following and maintains strong performance at very large context windows (reported up to 1M tokens). Considering the higher token cost, we choose the GPT-4.1-mini in evaluation. Gemini-2.0-Flash (DeepMind, 2025) targets high throughput and the use of built-in tools, offering a 1M token context window for efficient long-context processing. Claude-3.7-Sonnet (Anthropic, 2025) is a hybrid-reasoning model with optional visible “extended thinking,” strong math/coding skills, and developer-controlled thinking budgets.
C.2 RAG Agents
@@
We implement a BM25 (Robertson and Walker, 1994) retriever as a strong lexical baseline: it scores documents by term frequency with saturation and inverse document frequency, with length normalization controlled by parameters and . BM25 remains competitive for exact-match queries and complements dense retrievers with robust precision on keyworded questions.
@@
C.3 Agentic Memory Agents
For Agentic Memory Agents, We evaluate the Self-RAG (Asai et al., 2023), MemGPT (Packer et al., 2023), and MIRIX (Wang and Chen, 2025) on our benchmark. Self-RAG use LLMs as the agent to decide when/what to retrieve and to critique its own outputs. MemGPT operates the hierarchical memory management, paging relevant snippets between short-term and long-term stores and using event-driven interrupts to maintain coherence and evolvability over extended interactions. MIRIX adopts a multi-agent memory architecture with six specialized memory types (Core, Episodic, Semantic, Procedural, Resource, Knowledge Vault) and a coordinator that orchestrates updates/retrieval across agents.
@@ Appendix D Prompts
D.1 Instructions for Memory Construction
@@
D.2 Instructions for Task Execution
@@
Appendix E Detailed Experimental Results
@@
E.1 Detailed Results on TTL
We give detailed results on Multi-Class Classification (MCC) task in Table E. For all three types of tasks, RAG-based agents generally underperform compared to their respective GPT-4o-mini backbones. This observation highlights certain limitations inherent to the RAG approach. For instance, in TTL tasks, RAG-based methods often struggle to more accurately retrieve context from memory that is closely associated with the input.
E.2 Results on Input Chunk Size Ablation Study
@@
E.3 Results on Retrieval TopK Ablation Study
@@
E.4 Results on Different Context Length Ablation Study
@@
E.5 Results on Computational Latency Analysis
To illustrate the latency of various memory agents in terms of (1) Memory Construction (M.C.); (2) Query Execution (Q.E.), we report the latency of various memory agents on MH-QA and LME (S*). This part of experiments is done on a server with Four NVDIA L40 GPU and AMD EPYC 7713 64-Core CPU. We use the NV-Embed-v2 (7B) as the embedding model in HippoRAG-v2. We show the results in Table 12 and 12. From the table, we find that using a smaller chunk size requires significantly more time for memory construction, especially for methods such as HippoRAG-v2, Mem0, Cognee, and MemGPT. Meanwhile, methods such as Mem0, Cognee and MIRIX need extremely high resources when constructing the memory.
@@ E.5 Results on Computational Latency Analysis
E.6 GPU Memory Usage Comparison
@@
Appendix F Experimental Settings
@@ F.2 Settings of the RAG Agents
F.3 Settings of the Chunk Size
@@
Appendix G Task Rationale and Justification for Selective Forgetting Task
While the Selective Forgetting task may appear specialized or even synthetic at first glance, it is designed to address a fundamental, universal challenge in long-term memory systems: maintaining context efficiency and mitigating interference between outdated and updated information. We justify the design, novelty, and validity of this task along four tightly linked dimensions:
1.
Theoretical Necessity: For any memory system—biological or artificial—storage capacity is inherently finite. The ability to autonomously discard (or selectively forget) outdated, redundant, or superseded information is not a niche requirement, but a core prerequisite for keeping memory representations concise, robust, and free from conflicting signals. Our task is designed as a controlled proxy to evaluate this underexplored yet critical capability.
2.
Distinction from Previous Settings (e.g., Knowledge Updating): While prior work has explored knowledge updating (which focuses on overwriting old facts with new, conflicting ones), our work uniquely emphasizes the explicit, proactive removal of non-essential information to free up cognitive and contextual space. This distinguishes our task from existing fact-updating benchmarks, and we view this framework as a foundational step toward evaluating more complex memory management behaviors in real-world settings.
3.
Justification for the Controlled Synthetic Setting: We acknowledge that the current task setup includes synthetic elements, and this design choice is deliberate and methodologically justified. Constructing a naturalistic dataset with long-term interaction history (over 100K tokens) and unambiguous, precise ground-truth annotations for what information should be forgotten is inherently challenging, as real-world forgetting decisions are often ambiguous and context-dependent. Our controlled synthetic setting isolates the selective forgetting capability from confounding factors, enabling reproducible, apples-to-apples comparisons across different memory agent architectures.
4.
Validity and Feasibility of the Task: Critically, we confirm that our proposed task, despite its synthetic design, is fully valid and solvable. As demonstrated in our ablation study, long-context agents with strong reasoning capabilities achieve near-perfect performance on short-context (6K tokens) versions of the dataset. This result directly confirms that performance degradation on long-context, full-length task inputs stems not from flaws in the task definition itself, but from the fundamental limitations of current memory agents in long-range reasoning—specifically, their inability to accurately identify and discard outdated information across extended interaction histories.
Appendix H Supplementary Validation for Test-Time Learning
This section provides supplementary justification for the terminology and design of our Test-Time Learning (TTL) evaluation paradigm, as well as additional zero-shot baseline experiments that validate the core premise of the TTL task.
H.1 Terminology and Design Rationale
We use the term Test-Time Learning (TTL) to describe the class of tasks that evaluate an agent’s ability to acquire task-specific skills and rules incrementally from interaction history, and apply these learned patterns to unseen inputs at inference time. The terminology and task design are justified by three core principles:
1.
Distinguishing Skill Acquisition from Static Information Retrieval: We explicitly differentiate TTL from standard retrieval-focused tasks (e.g., Accurate Retrieval, Long-Range Understanding) to highlight its unique focus on learning, rather than fact recall. In retrieval tasks, the agent’s core goal is to fetch pre-defined, static facts from the interaction history. In contrast, TTL tasks (e.g., multi-class classification on Banking77, personalized movie recommendation) require the agent to induce latent classification rules, preference patterns, and task schemas from sequential labeled examples in the dialogue history, then generalize these patterns to completely novel, out-of-sample inputs. This operationalization directly aligns with the canonical definition of learning: updating a behavioral policy based on accumulated experience, which occurs exclusively during test-time interaction with no pre-training or fine-tuning on the target task.
2.
Controlled Operationalization of Online Learning: While a fully dynamic online learning setting would involve an interleaved loop of memory update, retrieval, task execution, and feedback-driven memory refinement, such a setup introduces significant confounding variables that make it impossible to isolate an agent’s memory capacity from unrelated reasoning, planning, or execution errors. To enable robust, reproducible, and unbiased evaluation of pure memory-enabled learning capability, we adopt a two-stage protocol that retains the core of online learning while eliminating confounds:
Acquisition Phase: The agent incrementally processes a sequence of labeled task examples, simulating the experience accumulation process of real-world agent interactions over time.
Evaluation Phase: The agent is tested on its ability to apply the rules and patterns acquired in the acquisition phase to held-out, unseen inputs.
This design ensures that performance differences between models can be directly attributed to their ability to leverage long-term interaction history for learning, rather than spurious factors from dynamic feedback loops.
3.
Foundational Framework for Agent Memory Research: As research on memory-enabled LLM agents remains in a nascent stage, our TTL evaluation framework provides a standardized, reproducible operationalization of in-situ learning for memory agents. While our current protocol simplifies the fully online learning loop for evaluation stability, it successfully captures the core competence required for real-world self-improving agents: the ability to improve task performance solely by memorizing and generalizing from past interactions. We will extend this framework to more complex, fully interleaved online learning scenarios in future work.
H.2 Zero-Shot Baseline Validation Experiments
A core premise of our TTL task design is that performance improvements in the full memory setting are driven by the agent’s ability to learn from historical examples, rather than prior knowledge encoded in the base LLM’s pre-training. To validate this premise, we conducted zero-shot baseline evaluations, where models were tested on the TTL tasks with no access to the historical example sequence (i.e., no opportunity for test-time learning).
Table 16 presents the zero-shot performance of three mainstream LLMs on our two core TTL tasks: Multi-Class Classification (MCC, Banking77) and personalized Movie Recommendation (Recom.). We also include the full memory performance of GPT-4o-mini for direct comparison, to quantify the performance degradation when test-time learning is disabled.
| Model | MCC | Recom. | Avg. |
| Zero-Shot Setting (No Access to Historical Examples) | |||
| GPT-4o-mini | 0.6 | 6.1 | 3.4 |
| GPT-4.1-mini | 0.8 | 5.7 | 3.3 |
| Gemini-2.0-Flash | 0.0 | 5.5 | 2.8 |
| Full Memory Setting (With Access to Historical Examples) | |||
| GPT-4o-mini w/ full context | 82.0 | 15.1 | 48.6 |
The results confirm our core hypothesis: all models exhibit near-chance performance in the zero-shot setting, with average accuracy below 4% across both tasks. In contrast, GPT-4o-mini achieves 48.6% average accuracy when provided with the full historical example sequence, representing a 45.2 percentage point absolute improvement. This stark performance gap demonstrates that the base LLMs have no meaningful prior knowledge to solve these long-tail tasks out of the box, and all performance gains in the full memory setting are explicitly driven by the agent’s ability to learn from the provided interaction history. This validates that our TTL benchmark accurately measures test-time learning capability, rather than pre-training knowledge or spurious pattern matching.
Appendix I Cost-Performance Analysis
I.1 Methodology
To provide a realistic assessment of the practical limitations of each agent architecture, our cost calculation assumes Context Caching is enabled (a standard feature in modern APIs like OpenAI), which significantly reduces the cost for Long-Context (LC) models processing shared histories. We compare three representative architectures: Long-Context (LC) models, RAG Agents, and the agentic memory system MIRIX.
Pricing Basis: Costs are calculated based on OpenAI’s pricing as of November 2025:
GPT-4o-mini: $0.15 / 1M input tokens, $0.60 / 1M output tokens
GPT-4.1-mini: $0.40 / 1M input tokens, $1.60 / 1M output tokens
Context Caching: We apply cached input pricing (50% discount for GPT-4o-mini, 75% discount for GPT-4.1-mini) for Long-Context Agents, assuming sequential questioning on shared contexts.
Settings: RAG agents use Top-K=10. Embedding indexing costs are excluded as one-time expenses.
I.2 Cost-Performance Results
We report the amortized inference cost per query (in USD) and the corresponding performance metric (Accuracy/Score) across four representative datasets that vary in context length and reasoning complexity in Table 17.
| Model/Architecture | MH-Doc QA | MCC | Detective QA | FC-SH | ||||
| Est. Cost (USD) | Performance | Est. Cost (USD) | Performance | Est. Cost (USD) | Performance | Est. Cost (USD) | Performance | |
| GPT-4o-mini | $0.01 | 43.0 | $0.008 | 82.0 | $0.01 | 63.4 | $0.01 | 45.0 |
| GPT-4.1-mini | $0.043 | 66.0 | $0.011 | 75.6 | $0.013 | 56.3 | $0.027 | 36.0 |
| RAG Agents (BM25 + 4o-mini) | <$0.001 | 56.0 | $0.006 | 75.4 | $0.006 | 52.1 | <$0.001 | 48.0 |
| MIRIX (4.1-mini) | $0.016 | 75.0 | $0.010 | 61.0 | $0.011 | 62.0 | $0.019 | 20.0 |
I.3 Key Insights
1.
The Efficiency-Reasoning Trade-off in RAG: RAG agents like BM25 are cost-efficient (< $0.001 – $0.006 per query) but suffer in tasks requiring global reasoning (e.g., Detective QA, scoring 52.1 vs. 63.4 for LC models). This limits their utility in complex analytical scenarios.
2.
Cost Prohibitiveness of Long-Context Scaling: While powerful, Long-Context models face steep cost increases with stronger backbones. Upgrading from GPT-4o-mini to GPT-4.1-mini roughly quadruples the cost (from $0.010 to $0.043 for MH-Doc QA), making high-end long-context deployment expensive even with caching.
3.
Agentic Memory (MIRIX) as the Optimal Middle Ground: A critical finding is that MIRIX (with GPT-4.1-mini) achieves a lower cost ($0.016) than the raw GPT-4.1-mini Long-Context setup ($0.043) on memory-intensive tasks like MH-Doc QA, while delivering superior performance (75.0 vs. 66.0). This demonstrates that agentic memory mechanisms can successfully decouple performance from linear context costs, offering a scalable solution for high-performance applications.
Appendix J Strict Compute-Matched Comparative Experiments
J.1 Experimental Setup
To address concerns about budget effects confounding the comparison between architectures, we conducted a strict compute-matched ablation study on Banking77 (TTL) and Book Summarization (LRU) using the strongest backbone (GPT-4.1-mini). We define three budget levels:
Low ( 4K tokens): Constrain Long-Context (LC) models to truncated contexts; limit RAG and MIRIX to Top-K=1 chunks to match token counts.
Medium ( 40K tokens): Moderate budget setting for all architectures; limit RAG and MIRIX to Top-K=10 chunks.
High ( 100K+ tokens): Scale RAG and MIRIX to retrieve high Top-K chunks to match the LC full-context budget.
For Book Summarization, we evaluate on a random 30-book subset for computational efficiency. We compare total compute load via total processed tokens, rather than strictly equalizing forward passes, as forcing agentic models to a single forward pass would strip them of their core reasoning capabilities.
| Task | Budget Setting | Long-Context (GPT-4.1-mini) | RAG (BM25) | Agentic (MIRIX) |
| TTL (Banking77) | Low ( 4K / Top-K=1) | 74.0 | 83.0 | 52.0 |
| Medium ( 40K / Top-K=10) | 90.0 | 89.0 | 65.0 | |
| High ( 104K) | 93.0 | 88.0 | 67.0 | |
| LRU (Book Summarization) | Low ( 4K / Top-K=1) | 8.2 | 7.9 | 8.4 |
| Medium ( 40K / Top-K=10) | 16.4 | 15.8 | 18.7 | |
| High ( 113K) | 39.7 | 38.0 | 38.8 |
We summarize our key findings as follows:
TTL: The Efficiency-Capacity Trade-off: At the Low ( 4K) budget, RAG (83.0) significantly outperforms LC models (74.0), confirming RAG’s superior structural efficiency for pattern matching via precise retrieval of relevant examples, while LC models suffer heavily from truncation. As the budget increases to Medium ( 40K), performance equalizes ( 90.0). At the High budget, LC models scale to 93.0, whereas RAG saturates and slightly degrades (88.0) due to retrieved noise. This demonstrates that while RAG is efficient, LC models have a higher capacity ceiling when the budget permits.
LRU: The Information Threshold Effect: For global reasoning tasks, performance exhibits a clear threshold behavior. At Low and Medium budgets, all architectures fail (<20.0 score), confirming that partial information is insufficient for full-book summarization regardless of the method. Only at the High budget (full text access) do all models achieve meaningful performance ( 39.0), with RAG nearly matching LC models. This proves that for LRU tasks, success is determined by meeting the full information threshold, not by the architecture itself.
Appendix K Prompt Design and Overwrite Policy Ablation Experiments
K.1 Prompt Design Specification
Unlike standard long-context evaluations that input raw text, we wrap all input chunks within a simulated User-Assistant dialogue to explicitly trigger the agent’s memory mechanism. Each input chunk is preceded by a memorization instruction (e.g., "Please memorize the following information for future questions") to establish a clear intent for information storage. For each specific dataset, we carefully curated task instructions to ensure agents accurately comprehend the task intent and execute the required actions.
Crucially, for the Selective Forgetting competency, we introduce explicit guardrails in the prompt. We explicitly instruct agents that facts are indexed by serial numbers, and that newer facts have larger serial numbers. Agents are mandated to resolve conflicts by prioritizing the newest fact (full prompt templates are provided in the supplementary code repository).
We clarify that while prompt deltas can shift outcomes, our benchmark applies a unified and standardized prompt template across all evaluated agents (Long-context, RAG, and Agentic). This ensures that the performance gaps observed (e.g., the failure of RAG agents in multi-hop Selective Forgetting) are attributed to the limitations of their memory mechanisms rather than prompt inconsistency.
K.2 Overwrite Policy Ablation Experiments
To rigorously test whether explicit instructions can mitigate the forgetting/overwriting issue, we conducted additional ablation studies using explicit overwrite prompts on the GPT-4.1-mini baseline. We test two policy settings:
Policy A (Always Prefer Later): "Crucial Rule: Treat the facts as a chronological update stream. If there is ANY conflict between facts, you must ALWAYS overwrite the earlier fact with the one having the larger serial number."
Policy B (Conservative/Explicit Negation): "Crucial Rule: Be conservative with updates. ONLY discard or overwrite an earlier fact if the fact with the larger serial number explicitly negates it or explicitly states the previous information is incorrect."
K.2.1 Ablation Results
The results of the overwrite policy ablation are presented in Table 19.
| Model/Setting | FC-SH | FC-MH | Avg. |
| GPT-4.1-mini (Baseline) | 36.0 | 5.0 | 20.5 |
| GPT-4.1-mini (Policy A) | 40.0 | 4.0 | 22.0 (+1.5) |
| GPT-4.1-mini (Policy B) | 28.0 | 4.0 | 16.0 |
K.2.2 Key Insights
1.
Limited Generalization with Aggressive Updates: While Policy A slightly improves performance on single-hop tasks (FC-SH score increases from 36.0 to 40.0), it fails to generalize to complex multi-hop reasoning (FC-MH drops to 4.0). This suggests that while prompting helps with simple fact retrieval, it cannot effectively propagate updates through multi-step reasoning chains.
2.
Performance Degradation with Conservative Constraints: Policy B significantly degrades average performance (-4.5 points). The complex conditional instruction (checking for explicit negation) increases the model’s cognitive load and induces overly cautious behavior, preventing the retrieval of valid updates.
These findings serve as a sanity check that validates our core motivation: Selective Forgetting cannot be solved by prompt engineering alone, and requires dedicated memory mechanism design.
Appendix L Supplementary Notes for LLM-as-a-Judge and Input Format
L.1 Validity of LLM-as-a-Judge Evaluation
We acknowledge that model-based evaluation may not be optimal for all task contexts. However, we adopted LLM-as-a-judge scoring on the LongMemEval and Bench-Sum datasets to remain consistent with prior work, and we validate its appropriateness as follows:
For LongMemEval, the questions admit clear, objective ground-truth answers with minimal subjectivity. Wu et al. (2025) report that a prompt-engineered GPT-4o judge achieves 98.0% agreement with human annotations, demonstrating very high stability.
Yen et al. (2024) validate that GPT-4o judgments mostly align with human evaluations for long-context summarization tasks.
These findings confirm that our LLM-as-a-judge setup reliably reflects human evaluation for the tasks in our benchmark.
L.2 Rationale for Chunked Input Format
We acknowledge that real-world personal assistants may receive streaming input (e.g., continuous user interactions or real-time data streams). However, in practice, it is necessary to quantize real-world inputs before feeding them into a language model (e.g., accumulating continuous input into discrete chunks for inference). Thus, inputting chunks into the agent is a natural and realistic strategy to handle streaming input in real-world deployments.
Additionally, the chunked input format simulates the incremental, multi-turn nature of real-world user-agent interactions, where information arrives sequentially over time rather than in a single full document. This setup is critical for evaluating memory agents, which are designed to process information incrementally, rather than the static full-document input used in standard long-context benchmarks.